# Refinement-based Christoffel sampling for least squares approximation in non-orthogonal bases

### v1

@@ 5 Numerical experiments
    Extreme learning machines and neural networks. We consider a single-hidden-layer-feedforward neural network with fixed, randomly initialized hidden parameters. Such networks are often referred to as extreme learning machines (ELMs) [13]. When considering additive nodes, the basis equals
    (26) $\Phi_{n}=\{g(a_{i}x+b_{i})\}_{i=1}^{n}\qquad$
    where $a_{i}\in\mathbb{R}^{d},b_{i}\in\mathbb{R}$ are randomly intialized and $g$ is a bounded nonconstant piecewise continous activation function. Universal approximation properties for such networks on compact input sets are proven in [26].
    We consider (26) on $0,1]^{2}$ with $a_{i}$ uniformly randomly sampled in $1,1]^{2}$ and $b_{i}$ uniformly randomly sampled in $[0,1]$, $g$ sigmoidal $\left(g(x)=\frac{1}{1+\exp(-x)}\right)$ and $n=600$. Figure 8 depicts the first two basis function associated with this ELM and the output of the RCS algorithm.
    Note that similar techniques can be applied even when the hidden nodes are not fixed. Specifically, the output of a deep neural network can be viewed as a linear combination of the outputs of basis functions associated with the nodes in the penultimate layer. This perspective, often called the adaptive basis viewpoint, naturally leads to hybrid training procedures: the weights in the last layer are learned via least squares, while all other weights are optimized using gradient descent [12]. In [3], the least squares component is further enhanced using Christoffel sampling.
    (a) $\phi_{1}$
    (b) $\phi_{2}$
    (c) $u$
    (d) sample points
    Figure 8: First basis functions and output of the RCS algorithm for one instance of an ELM (26) on $X=[0,1]^{2}$, with sigmoidal activation function, $n=600$, $a_{i}$ uniformly randomly sampled in $1,1]^{2}$ and $b_{i}$ uniformly randomly sampled in $[0,1]$.
Appendix A Discrete result
In this appendix, we prove a discrete result very similar to [10, Theorem 2] (for leverage scores) and [11, Lemma 6.2] (for ridge leverage scores with variable regularization parameter). The main difference is that the following result holds for ridge leverage scores with a fixed regularization parameter, which does not pose any new challenges compared to known results. We will use the following notation:
$\tau_{k}^{\epsilon}(A)$ $=a_{k}^{\top}\left(A^{*}A+\epsilon^{2}I\right)^{-1}(a_{k}^{\top})^{*}$
$n^{\epsilon}(A)$ $=\sum_{k=1}^{m}\tau_{k}^{\epsilon}(A)=\sum_{i=1}^{n}\frac{\lambda_{i}(A^{*}A)}{\lambda_{i}(A^{*}A)+\epsilon^{2}}$
where $a_{k}^{\top}$ is the $k$-th row of the matrix $A$. This Lemma is used to prove Lemma 4.3.
Lemma A.1.
Let $A\in\mathbb{C}^{m\times n}$, $u\in\mathbb{R}^{m}$ with $u_{k}>0$ for all $1\leq k\leq m$ and $\epsilon>0$, then there exists a diagonal matrix $W\in\mathbb{R}^{m\times m}$ with $0\preceq W\preceq I$ such that
(27) $\tau_{k}^{\epsilon}(WA)\leq u_{k},\qquad 1\leq k\leq m,$
and
(28) $\sum_{k:W_{kk}\neq 1}u_{k}\leq n^{\epsilon}(A).$
Proof A.2.
The proof is very similar to that of [10, Theorem 2]. Note that for $\epsilon>0$, every ridge leverage score is strictly smaller than one since $(WA)^{*}WA+\epsilon^{2}I\succ(WA)^{*}WA$ and, hence,
(29) $\tau_{k}^{\epsilon}(WA)<\tau_{k}(WA)\leq 1.$
The “whack-a-mole” algorithm can, therefore, be adapted to Algorithm 2. Similarly, for all $j\geq 0$, we let $W^{(j)}$ be the value of the $j$-th update to the weight and we show that $\overline{W}=\lim_{j\to\infty}W^{(j)}$ meets the conditions. Using the same arguments as in the proof of [10, Theorem 2], it follows from [10, Lemma 5] (which can easily be shown to hold also for $\tau_{k}^{\epsilon}$) that $W^{(j)}$ exists, its entries are nonnegative for all $j\geq 0$ and $\overline{W}$ exists. Furthermore, since $\epsilon>0$ and $W^{(j)}$ is nonnegative, it is easy to see that $\tau_{k}^{\epsilon}(WA)$ is continuous in $W$. Hence, by construction we have
$\tau_{k}^{\epsilon}(\overline{W}A)=\lim_{j\to\infty}\tau_{k}^{\epsilon}(W^{(j)}A)\leq u_{k},\qquad\forall k.$
It only remains to show that (28) holds for $W=\overline{W}$. Let $j$ be the first iteration such that $W_{kk}^{(j)}\neq 1$ for any $k$ such that $\overline{W}_{kk}\neq 1$. Following [10, Lemma 5], decreasing the weight of one row can only increase the ridge leverage scores of all other rows. This implies that
$\sum_{k:\overline{W}_{kk}\neq 1}u_{k}$ $\leq\sum_{k:\overline{W}_{kk}\neq 1}\tau_{k}^{\epsilon}(W^{(j)}A)$
$\leq n^{\epsilon}(W^{(j)}A)$
$=\sum_{i=1}^{n}\frac{\lambda_{i}((W^{(j)}A)^{*}(W^{(j)}A))}{\lambda_{i}((W^{(j)}A)^{*}(W^{(j)}A))+\epsilon^{2}}$
$\leq\sum_{i=1}^{n}\frac{\lambda_{i}(A^{*}A)}{\lambda_{i}(A^{*}A)+\epsilon^{2}}$
$=n^{\epsilon}(A),$
where in the penultimate step we use that $(W^{(j)}A)^{*}W^{(j)}A\preceq A^{*}A$ and, hence, $\lambda_{i}((W^{(j)}A)^{*}W^{(j)}A)\leq\lambda_{i}(A^{*}A),\forall i$ [25, §9-5], as well as the fact that $\lambda/(\lambda+\epsilon^{2})$ is monotonically increasing for nonnegative $\lambda$.
Algorithm 2 Whack-a-mole algorithm
1:Initialize $W=I$
2:while true do
3:  for $k=1$ to $m$ do
4:   if $\tau_{k}^{\epsilon}(WA)\geq u_{k}$ then
5:      Decrease $W_{kk}$ so that $\tau_{k}^{\epsilon}(WA)=u_{k}$.
6:   end if
7:  end for
8:end while
9:return $W$
Appendix B Auxiliary results
Lemma B.1.
Let $X\subseteq\mathbb{R}^{d}$, $\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$, $\epsilon>0$ and $v:X\to\mathbb{R}$ be a function satisfying
$0\leq v(x)\leq 1,\qquad\forall x\in X.$
Then,
$n^{\epsilon}(v\Phi)\leq n^{\epsilon}(\Phi).$
Proof B.2.
Denote the gram matrices with respect to the $L^{2}_{\rho}(X)$ inner product associated with the sets $\Phi$ and $v\Phi$ by $G$ and $G^{v}$, respectively. Since
$(G)_{ij}-(G^{v})_{ij}=\int_{X}(1-v(x)^{2})\phi_{i}(x)\phi_{j}(x)d\rho=\langle\sqrt{1-v^{2}}\phi_{i},\sqrt{1-v^{2}}\phi_{j}\rangle_{L^{2}_{\rho}(X)}$
one has $G-G^{v}\succeq 0\Rightarrow G\succeq G^{v}$. This implies that $\lambda_{i}(G^{v})\leq\lambda_{i}(G),\forall i$ [25, §9-5], which allows us to bound
$n^{\epsilon}(v\Phi)=\sum_{i=1}^{n}\frac{\lambda_{i}(G^{v})}{\lambda_{i}(G^{v})+\epsilon^{2}}\leq\sum_{i=1}^{n}\frac{\lambda_{i}(G)}{\lambda_{i}(G)+\epsilon^{2}}=n^{\epsilon}(\Phi),$
using the fact that $\lambda/(\lambda+\epsilon^{2})$ is monotonically increasing for nonnegative $\lambda$.
Lemma B.3.
Let $X\subseteq\mathbb{R}^{d}$, $\{\phi_{i}\}_{i=1}^{n}\subset L^{2}_{\rho}(X)$, $M\in\mathbb{C}^{n\times n}$ positive semidefinite, $\epsilon>0$ and $\beta\geq 1$. Then,
$k^{\epsilon}(x;\Phi,M)\leq\beta k^{\epsilon}(x;\Phi,\beta M).$
Proof B.4.
One has
$k^{\epsilon}(x;v\Phi,M)=\phi(x)^{*}(G+\epsilon^{2}I)^{-1}\phi(x)=\beta\phi(x)^{*}(\beta G+\beta\epsilon^{2}I)^{-1}\phi(x).$
Furthermore, $\beta G+\beta\epsilon^{2}I\succeq\beta G+\epsilon^{2}I$ since $\beta\geq 1$ and, hence,
$k^{\epsilon}(x;v\Phi,M)\leq\beta\phi(x)^{*}(\beta G+\epsilon^{2}I)^{-1}\phi(x)=\beta k^{\epsilon}(x;\Phi,\beta M)$
as required.
@@ Acknowledgments
The authors wish to thank Daan Huybrechs and Christopher Musco for fruitful discussions on the topic.

### v2

@@ 5 Numerical experiments
    Spectral method for fractional Laplacians. In [papadopoulos2023frame], a spectral method is developed for fractional differential equations. More specifically, an approximation $\hat{v}$ to the solution vv—that decays as $\lvert x\rvert\to$$\infty$—of an equation $\mathcal{L}[v]=f$ on $\mathbb{R}^{d}$ involving the fractional Laplacian $(-\Delta)^{s}$, $s\in(0,1)$, is sought in the span of extended and weighted orthogonal polynomials $\{\phi_{i}\}_{i}$,
    $v\approx\hat{v}=\sum_{i}c_{i}\phi_{i},\qquad\mathcal{L}\hat{v}=\sum_{i}c_{i}\,\mathcal{L}\phi_{i}\approx f.$
    This equation can be discretized by sampling, a procedure known as collocation. Hence, finding effective collocation points is equivalent to identifying good sampling points for approximation in the span of $\{\mathcal{L}\phi_{i}\}_{i}$, which can be achieved using the RCS algorithm. To illustrate this, we consider the two-dimensional setting described in [papadopoulos2023frame, §7.4], restricting the domain to the disk of radius 10 as proposed in that example. For further details, we refer to [papadopoulos2023frame, §7.4]. Figures 2(d) and 4(d) show the function $u$ and the sample points computed by the RCS algorithm, while Figure 3(d) presents the convergence results comparing uniform samples, RCS samples (using $C_{3}=5$) and the deterministic samples proposed in the original paper. Note that the error plot reports $\|v-\hat{v}\|_{\infty}$ on the disk of radius 10, which differs from $\|f-\mathcal{L}\hat{v}\|_{\infty}$. We find that RCS gives the same level of accuracy, while using far fewer samples than the deterministic procedure.
(a) polynomials + weighted polynomials
(b.1) lightning approximation
(b.2) 2D lightning approximation
(c) Fourier extension
(d) fractional spectral method
Figure 2: Illustration of the function $u$ computed by the RCS algorithm. Optionally, a comparison is made with an approximation to $k_{n}^{\epsilon}$ obtained via the dense grid method [dolbeault2022optimal, adcock2020near, migliorati2021multivariate] (dashed line) using a sufficiently dense grid. For more details, see the subsections in §5 corresponding to each plot label.
(a) polynomials + weighted polynomials
(b.1) lightning approximation
(b.2) 2D lightning approximation
(c) Fourier extension
(d) fractional spectral method
Figure 3: Convergence of discrete least squares approximation using samples computed by the RCS algorithm (circles) and the same number of uniformly random samples (triangles). The line indicates the geometric mean and the shaded area shows the variance over ten repetitions. The reported error is computed on an independent, dense grid. In (d), the deterministic sample strategy proposed in [papadopoulos2023frame, §7.4] is added to the comparison (crosses), including a comparison of the number of sample points. For more details, see the subsections in §5 corresponding to each plot label.
(b.2) 2D lightning approximation
(d) fractional spectral method
Figure 4: Illustration of the sample points outputted by the RCS algorithm. For more details, see the subsections in §5 corresponding to each plot label.
Figure 5: Comparison between the RCS algorithm (using $C_{3}=15$) and the discrete grid method [dolbeault2022optimal, adcock2020near, migliorati2021multivariate], which approximates $k^{\epsilon}_{n}$ by discrete orthogonalization of the basis on a uniformly random grid of $\ell$ points, for the 1D lightning approximation. Left: error evaluated on an independent, dense grid; right: elapsed time required to compute sample points. The experiment was conducted on a contemporary laptop using a Julia implementation [github_repo].
6 Conclusions
The refinement-based Christoffel sampling (RCS) algorithm provides an efficient and robust approach for discretizing non-orthogonal bases, yielding stable and accurate least-squares approximations. By iteratively refining a random sampling measure, the method naturally adapts to the underlying geometry of the basis while accounting for the effects of finite precision
There remain, however, several opportunities for further improvement. In the highly redundant regime, where the numerical dimension $n^{\epsilon}$ is much smaller than $n$, there still seems to be room for further computational savings. In particular, the evaluation of $u$ may be accelerated via randomized numerical linear algebra, since the matrix $A$ formed at each iteration is expected to be low rank and therefore suitable for oblivious dimensionality-reduction techniques [avron2017sharper]. Developing reliable and practical estimators for the numerical dimension—see also the discussion in §3.3—is another important step toward making the scheme effective in this setting. Finally, a more general extension would be to adapt the algorithm to unbounded domains, for which $k_{n}^{\epsilon}$ may fail to be uniformly bounded.
@@ Acknowledgments
The authors wish to thank Daan Huybrechs and Christopher Musco for fruitful discussions on the topic. BA was supported by the the Natural Sciences and Engineering Research Council of Canada (NSERC) through grant RGPIN-2021-611675. AH is a PhD fellow of the Research Foundation Flanders (FWO), funded by grant 11P2T24N, and was additionally supported by an FWO travel grant for a long stay abroad (V412525N).